OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

Equations?...

OpenStudy (anonymous):

y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

I've tried to get a grasp on the assessment through flvs i can't really get a good concept though there lesson.

OpenStudy (anonymous):

It's: y = -6x - 15 The formula is y = mx + b. They have given you the slope, which is m = -6 and the points x and y. Now you just have to plug in the slope and the points in y = mx + b and it should look like this: -3 = -6(-2) + b. Now find b by solving the equation above and you get -15 :) And voila, you have everything to make an equation from y = mx + b to y = -6x - 15 :) ...Sorry if I am not too good with explaining :(

OpenStudy (anonymous):

Nah it was good. A lot more short, sweet, and to the point then the lesson i read. it's so dang elaborate it's irritating.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Are you any good with finding perpendicular lines with equations? cause that's the worst one for me.

OpenStudy (anonymous):

Haha it's been like almost three years since I touched math but if I look it over/study it then I will be able to solve it. :) What's the questions?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

oops hold on

OpenStudy (anonymous):

]Choose the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3x - 7.

OpenStudy (anonymous):

The answer is y = -3x + 10 :) Okay, so from what I read: To find the slope m of the equation you are trying to find, you take the given equation y = 1/3x - 7 and take its slope; m = 1/3: Now the slope of an equations line that is perpendicular is the negative inverse: which makes m = 1/3 to m = -3 (1/3 becomes 3 and positive sign becomes negative, therefore, -3): Now you have the slope m = -3 and the point (4,-2): Next you would plug in these values into y = mx + b, which gives you -2 = -3(4) + b Now you would solve for b and get 10. Now you have the slope m = -3 and b = 10 now plug them into y = mx + b and get y = -3x + 10 :) ... I hope it was helpful! This is what I understood from what I read :)

OpenStudy (anonymous):

Now if you were given an equation in the form of y - y1 = m(x - x1), you would change it to the form y = mx + b and repeat the steps I wrote above :)

OpenStudy (anonymous):

oh okay see i think that's what confused me. i kept taking the point slope formula and trying to transition that to slope intercept. And i didn't realize all you had to do with the slope intercept formula is plug in the numbers.I Appreciate you're time.

OpenStudy (anonymous):

No problem :) I am here to help haha :)